# SpecMaskFoley: Steering Pretrained Spectral Masked Generative Transformer Toward Synchronized Video-to-audio Synthesis via ControlNet

### v1

@@ 3.1 SpecMaskGIT
SpecMaskGIT features a distillation-free approach to efficient high-quality TTA synthesis [3]. The efficiency, effectiveness and flexibility of this method are based on: (1) A highly compressed latent space created by a 2-D SpecVQGAN [30]; (2) the light-weight ViT backbone trained with the mask reconstruction task ([31, 32, 8]) upon the latent space of SpecVQGAN; (3) the use of parallel iterative synthesis, i.e., MaskGIT sampler, for fast inference ([25, 33]).
@@ 3.3 ControlNet and FT-Aligner
As shown in Fig.2, we implement a Transformer version of ControlNet similar to [35, 36, 37], which initializes the ControlNet branch with the pretrained model weights and then connects the ControlNet branch to the backbone via zero linear layers. Although this design was presented for continuous Diffusion Transformers [35, 36, 37], we found it works well in our discrete model. Note that the only training loss in SpecMaskFoley is cross entropy loss, the standard training target for discrete models ([25, 33, 3, 8]).
@@ 3.4 Multi Classifier-Free Guidance
where $\ell_{\mathrm{uncond}}$ denotes the logits gained from the audio backbone without the CLAP conditioning, $\ell_{\mathrm{video}}$ denotes logits earned by conditioning SpecMaskFoley with deep video features via ControlNet but without CLAP feautures in the backbone, $\ell_{\mathrm{text\&video}}$ denotes logits earned by conditioning SpecMaskFoley with both CLAP text features and deep video features simultaneously, and $t$ denotes the CFG scale. We do not use logits gained by conditioning the audio backbone with CLAP text features, and we also randomly replace the CLAP conditioning with an unconditonal mask (shown in Fig.2) for 90% of the training steps, as we found text prompts alone are less important in Sec.5.2.
@@ 4.1 Datasets
The pretraining of the TTA backbone in SpecMaskFoley was conducted on the summation of the unbalanced and balanced subset of AudioSet [43], a dataset that has been widely used in general audio representation learning ([44, 32]) and generative modeling ([2, 34]) due to its massive amount of audio clips as well as its diversity in sound sources and recording environments.
@@ 4.3 Metrics
We use the av-benchmark  [^1] https://github.com/hkchengrex/av-benchmark to evaluate the quality of foley synthesis from the following aspects: Audio synthesis quality. Following common practice ([1, 2, 3, 4, 6]), we compute the Frechet Distance (FD) and Kullback–Leibler (KL) distance on the basis of PaSST [44], a 32 kHz audio classifier. The 16 kHz VGGish classifier [47] is also used for FD (denoted as “FAD”). Note that we exclude PANN [48] from the metric computation, as it has been reported as not being robust in some scenarios  [^2] https://github.com/haoheliu/audioldm_eval. Audio-video semantic matching. The semantic similarity between the input video and the generated foley audio is evaluated by the cosine similarity between ImageBind [41] video and audio features, denoted as “IB similarity”. Audio-video synchronization. The synchronization between input video and generated audio is measured by DeSync [12] in seconds.
@@ 5.1 VGGSound Benchmarking
Last but not least, SpecMaskFoley remains competitive even compared with strong from-scratch baselines. methods. SpecMaskFoley substantially outperforms VATT [17], a from-scratch MaskGIT method, in most metrics in Tab. 1. Compared with V-AURA [18], SpecMaskFoley presents superior audio synthesis quality and inference speed, while maintaining the same level of audio-video synchronization. Overall, SpecMaskFoley performed similarly to Frieren [19], a from-scratch flow-matching method, but our method is more advantageous in metrics such as FAD, KL, IB similarity, and DeSync. Note that although SpecMaskFoley uses more parameters than Frieren and MMAudio, the trainable parameter amount is only 126M.
@@ 5.2 Ablation Studies and Discussions
Discussion: 2-D vs. 1-D VAE. As illustrated in Fig.2 and discussed in Sec.4.3, due to the use of a 2-D VAE in SpecMaskFoley, there are only 53 temporal frames in the latent space presenting a 10-second clip. However, the 1-D VAE used in Frieren and MMAudio presents a 10-second clip with more than 300 frames [2], preserving higher temporal resolution. According to the DeSync scores in Tab.1, this temporal resolution has not been an issue for SpecMaskFoley in the VGGSound benchmark. However, low temporal resolution may limit the potential of 2-D methods in the future. We leave the task of transplanting the success of SpecMaskGIT and SpecMaskFoley from 2-D to 1-D as our future work.

### v2

@@ 3.1 SpecMaskGIT
SpecMaskGIT features a distillation-free approach to efficient high-quality TTA synthesis [3]. The efficiency, effectiveness and flexibility of this method are based on: (1) A highly compressed latent space created by a 2-D SpecVQGAN [30]; (2) the light-weight ViT backbone trained with the Masked Language Modeling task ([31, 32, 8]) upon the latent space of SpecVQGAN; (3) the use of parallel iterative synthesis, i.e., MaskGIT sampler, for fast inference ([25, 33]).
@@ 3.3 ControlNet and FT-Aligner
As shown in Fig.2, we implement a Transformer version of ControlNet similar to [35, 36, 37], which initializes the ControlNet branch with the pretrained model weights and then connects the ControlNet branch to the backbone via zero-initialized linear layers. Although this design was presented for continuous Diffusion Transformers [35, 36, 37], we found it works well in our discrete model. Note that the only training loss in SpecMaskFoley is cross entropy loss, the standard training target for discrete models ([25, 33, 3, 8]).
@@ 3.4 Multi Classifier-Free Guidance
where $\ell_{\mathrm{uncond}}$ denotes the logits gained from the audio backbone without the CLAP conditioning, $\ell_{\mathrm{video}}$ denotes logits obtained from conditioning SpecMaskFoley with deep video features via ControlNet but without CLAP feautures in the backbone, $\ell_{\mathrm{text\&video}}$ denotes logits earned by conditioning SpecMaskFoley with both CLAP text features and deep video features simultaneously, and $t$ denotes the CFG scale. We do not use logits gained by conditioning the audio backbone with CLAP text features, and we also randomly replace the CLAP conditioning with an unconditonal mask (shown in Fig.2) for 90% of the training steps, as we found text prompts alone are less important in Sec.5.2.
@@ 4.1 Datasets
The pretraining of the TTA backbone in SpecMaskFoley was conducted on the sum of the unbalanced and balanced subset of AudioSet [43], a dataset that has been widely used in general audio representation learning ([44, 32]) and generative modeling ([2, 34]) due to its massive amount of audio clips as well as its diversity in sound sources and recording environments.
@@ 4.3 Metrics
We use the av-benchmark  [^1] https://github.com/hkchengrex/av-benchmark to evaluate the quality of foley synthesis from the following aspects: Audio synthesis quality. Following common practice ([1, 2, 3, 4, 6]), we compute the Frechet Distance (FD) and Kullback–Leibler (KL) distance on the basis of PaSST [44], a 32 kHz audio classifier. The 16 kHz VGGish classifier [47] is also used for FD (denoted as “FAD”). Note that we exclude PANN [48] from the metric computation, as it has been reported as not being robust in some scenarios  [^2] https://github.com/haoheliu/audioldm_eval. Audio-video semantic matching. The semantic similarity between the input video and the generated foley audio is evaluated by the cosine similarity between ImageBind [41] video and audio features, denoted as “IB similarity”. Audio-video synchronization. The synchronization between input video and generated audio is measured by DeSync ([6, 12]) in seconds.
@@ 5.1 VGGSound Benchmarking
Last but not least, SpecMaskFoley remains competitive even compared with strong from-scratch baselines. methods. SpecMaskFoley substantially outperforms VATT [17], a from-scratch MaskGIT method, in most metrics in Tab. 1. Compared with V-AURA [18], SpecMaskFoley presents superior audio synthesis quality and inference speed, while maintaining the same level of audio-video synchronization. Overall, SpecMaskFoley performed similarly to Frieren [19], a from-scratch flow-matching method, but our method is more advantageous in metrics such as FAD, KL, IB similarity, and DeSync. Note that although SpecMaskFoley uses more parameters than Frieren and MMAudio, the trainable parameter amount is only 126M, enabling single GPU training, hence is friendly to low-resource researchers.
@@ 5.2 Ablation Studies and Discussions
Discussion: 2-D vs. 1-D VAE. As illustrated in Fig.2 and discussed in Sec.4.3, due to the use of a 2-D VAE in SpecMaskFoley, there are only 53 temporal frames in the latent space presenting a 10-second clip. However, the 1-D VAE used in Frieren and MMAudio presents a 10-second clip with more than 300 frames [2], preserving higher temporal resolution. According to the DeSync scores in Tab.1, this low temporal resolution has not been a bottleneck for SpecMaskFoley in the VGGSound benchmark, but may limit the potential of 2-D methods in the future. We leave the task of transplanting SpecMaskGIT and SpecMaskFoley to 1-D as our future work.